On Tuesday 28th of February to Wednesday the 1st of March the OMX will be unavailable. A sCMOS PCO camera will be replaced and the system will be fully re-aligned.
The OMX has been serviced and new OTFs have been made. We recommend using these. A strange bug in the acquisition software causes the 405 nm laser line (1st channel) to present wrong values, i.e. intensity % of the laser in the drop-down list do not represent the actual values. The desktop background picture of the PC instructs you restore the right interface (in 3 clicks).
This week the filter drawer changer that was locked has been repaired. The system was serviced and softWoRx packages on the Windows master acquisition station and the workstations have been updated.
On the OMX imaging system, channel alignment is repeatedly disturbed by incorrect placement of the automated filter drawer (which sets in every time at startup) . A GE scientist is coming to solve this issue (in the week of February the 29th). Until then the filter drawer will be fixed to the BGR position to maintain correct aligning settings. We can work with it to image many but not all of your favourite dyes. Please take this into consideration.
The broken master PC on the OMX V4 system has been replaced and the OMX is fully functional again.
One of the PCs of the OMX-V4 system has a corrupted hard drive and does not boot. The system is not usable and will be repaired as soon as possible. The PC does not seem to be in stock in Europe, so it has been ordered in the US. As a result the system will not be operative for at least a week and likely for the rest if this month. We will do our best to speed up the process and will keep you posted in this blog.
The OMX V4 system is now updated with a DIC option to monitor ‘cell fitness’ or provide overviews of the cells that are imaged. So far, the quality is mediocre but we are trying to improve this.
The storage for gas tanks (CO2, N2) in the CCI corridor has been replaced and moved to make space for the flow cabinet that will be positionned in the recess behind it. Beware that the corridor has become more shallow because of the move.
A StorSimple backup system has been implemented by ITS. We have tested it for a couple of months and it seems to perform well. The extra network drive is available on all imaging machines (except for the Zeiss NLO system, that will be connected to the net in the coming weeks) and can be found at O:\Research\CCI\#equipment\##user##. Writing on the drive can only be done on our machines, users will be able to read only from these drives so data will be available but secured. We are implementing a secure logon system for users to read data from this drive.
Next week, we will be trained by GE healthcare again, now to improve alignment and calibration procedures as well as the DLM module to perform dSTORM type of measurements (alike the NIKON system). Unfortunately, only one of the 4 cameras can be used to perform DLM, Furthermore the system will be upgraded with DIC and the live cell option will be improved and tested.
A webcam has been installed to make finding regions on sample holders more easily.